Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Saturday, November 21, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 21, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/21/98 (Saturday)

Segment 1 report for 11/21/98: At Emidio Pump Station, electrical work continued at the substation, flooring was being installed in the office building. On Hwy. 166 excavation was completed to EMP .56. The bore operation continued at the California Aqueduct (EMP 12.28). At Grapevine Pump Station, electrical work and pipe fitting continued. On Tejon Bench, welders worked from EMP 16.67 to 16.83. Backfill and compaction continued to 17.17. Trench was excavated to EMP 18.10 (working south). On Grapevine Rd., welding took place from 16.08 to 15.98. Coating was completed to 16.16. Trench has been excavated to EMP 15.36. On the east side of the I-5/Grapevine bore, the coating crew completed the segment from 14.68 to 14.65. On the west side of the bore, shading and compaction continued from 14.64 to 14.49. A pull box was installed at EMP 14.58.

Segment 2B (Castaic to Sylmar): Report for 11-21-98. Drilling continued at the Santa Clara river while crews worked to contain an additional frac out. To aid in the removal of drilling mud from the frac outs an additional suction pump and tank were brought to the site. Meanwhile, trenching crews completed 500' of ditch to EMP 71.9 while 1000' of the road surface was cold planned to EMP 71.4. In addition, the Santa Susanna block valve was installed and backfilled near EMP 71.71. Coating and x-ray crews completed several joints to EMP 72.45 while bending and laying crews completed work at EMP 71.25. Near Sierra Hwy, 1000' of fiber optic duct was installed along with 500' of slurry to EMP 72.66.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 21 November 1998: 3A: Crews worked a late day to dig under two large storm drains north of Tujunga Wash today. Other work included the bore at Tuxford Street. This crew had equipment difficulties and did not achieve any significant footage. 3B: Fiber optic crews pulled two reels of cable in the MTA maintenance facility today. 3C: Concrete was poured for the cover at the high speed channel cover near the aqueduct crossing today.

Segment 4B(El Segundo to Lynwood): The crews at Lynwood Station and at Sepulveda ROW were filling the pipe whit water and preparing for the hydro test. The crew inside Lynwood station was working on the south wall. Segment 5 (Lynwood to Long Beach) Report for 21 November 1998: Boring operations continued slowly at Highway 91/Compton Creek. 38-inch swab was run. Drilling mud is being pumped into vacuum roll-away bins waiting for vacuum trucks. Bore at Wilmington Station is continuing.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page